# YORI: Autonomous Cooking System Utilizing a Modular Robotic Kitchen and a Dual-Arm Proprioceptive Manipulator

### v3

@@ II-A Dual-Arm Manipulator
Cobots emerged in the late 1990s, enabling safe human-robot collaboration in tasks such as organizing, assembly and inspection. While lighter and more responsive than traditional industrial robots, cobots still suffer from relatively slow acceleration and high stiffness that limit performance in dynamic environments, and their weight remains inefficient for mobile platforms. These limitations have driven research toward lighter, backdrivable, and compliant manipulators with improved force control, exemplified by Berkeley’s Quasi-Direct Drive (QDD) BLUE [1] and KoreaTech’s tendon-driven LIMS [2].
@@ II-B Proprioceptive Actuator
To enable backdrivability and compliance in manipulators, proprioceptive actuators, also known as QDD, offer an efficient solution. These actuators combine high-torque brushless DC motors with low reduction gearboxes (typically ¡10:1) to maintain torque transparency [3]. Unlike traditional high-geared servo motors that suffer from high reflected inertia and poor compliance, proprioceptive actuators excel at backdrivability and impact mitigation [4]. Their torque transparency enables direct torque estimation from motor current, eliminating the need for external force/torque sensors and simplifying both hardware and control systems.
@@ II-C Robotic Cell Layout Design
Robotic cell layout design extends traditional manufacturing layout optimization principles while addressing unique constraints of robotic systems, particularly in multi-robot environments and human-robot interaction scenarios [5]. Recent advances in computational power have enabled sophisticated optimization techniques including Mixed Integer Linear Programming, Genetic Algorithms, and Convex Optimization to solve complex workspace design problems, with simulation-based optimization and digital twin technologies allowing virtual validation before physical implementation.
@@ II-D Manipulator Control
Accurate modeling is crucial for controlling manipulators with proprioceptive actuators, enabling precise torque calculation and force feedback for stable operation in dynamic environments. This accurate modeling enables two key control strategies: inverse dynamics control for trajectory tracking [6] and impedance control for safe interaction tasks [7]
@@ II-E Motion Planning
Minimizing jerk (time derivative of acceleration) is essential for smooth trajectory generation. Flash and Hogan showed human arm movements naturally minimize jerk using $L_{2}$ norm optimization [8], while Yazdani demonstrated that $L_{\infty}$ norm minimization yields biologically plausible ”bang-bang” control policies [9]. Although sampling-based methods with time parameterization are computationally efficient in confined environments, they often sacrifice smoothness and energy efficiency, leading to continued research in optimization-based techniques with jerk constraints [10]. Our work leverages YORI’s structured kitchen environment to eliminate dynamic obstacles, enabling full exploitation of minimum jerk optimization for smooth, energy-efficient dual-arm trajectories.
@@ II-F Task Scheduling
Effective task scheduling and real-time adaptability are crucial for robust automated cooking systems. The kitchen can be modeled as a finite state machine of heterogeneous robotic appliances, with orders represented as finite sequences containing operational parameters like cooking duration and seasoning amounts. These constraints, formulated using linear temporal logic (LTL) and computation tree logic (CTL), naturally align with job-shop scheduling problems (JSSPs). Prior work has applied dispatching rules [11] and online MIP formulations [12] to cooking JSSPs. Building on these foundations, our work incorporates dynamic rescheduling logic and adaptive programming strategies to better handle real-world complexities including delays, failures, and continuous order arrivals.
@@ III Hardware Design
During the early stages of the project, we developed a prototype dual-arm manipulator using servo actuators to verify the feasibility of performing cooking tasks using robotic manipulators [13]. However, due to the usage of rigid servo actuators and lack of force/torque sensors, the robot manipulator was unable to complete certain tasks where a precise torque estimation or rapid movement was required, such as cutting or tossing food. Furthermore, since all the components including actuators and structural parts are off-the-shelf products without any design optimizations, the payload at the end-effector was limited and the overall workspace was not ideal. With the numerous constructive experiences from the servo motor actuated manipulator prototype, we developed the next-generation proprioceptive dual-arm manipulator using high-performance quasi-direct drive actuators. These actuators allow for the manipulators compatibility with many cooking tasks, in particular those requiring impacts and precise force control, while being compliant to changes in the constructed environment.
@@ III-A Dual-Arm Manipulator
An initial servo-actuated prototype validated the joint configuration and link dimensions for cooking tasks [13]. Building on this verified kinematic design, a next-generation dual-arm manipulator was developed using proprioceptive actuators, enabling impact-resilient manipulation with precise force control and environmental compliance, which are capabilities essential for dynamic cooking operations such as cutting and tossing.
@@ III-A1 Hardware Configuration
Link dimensions were determined based on human upper body proportions [14] for compatibility with existing kitchen tools: upper arm (0.400 m), forearm (0.375 m), and clavicle (0.420 m) lengths allow both coordinated dual-arm tasks and independent operation. Carbon fiber tubes were selected for all arm links due to their optimal strength-to-weight ratio. The 1.3 m platform height ensures full workspace coverage. Table II summarizes the mechanical parameters.
@@ III-A2 Actuator Selection
Joint torque requirements were determined through static moment analysis for worst-case loading conditions (fully extended horizontal configuration) [13]. Based on cooking task analysis, a 3 kg payload capacity was specified with a 1.5 safety factor, yielding maximum torque requirements of 37.93 Nm (shoulder pitch), 20.29 Nm (elbow pitch), and 3.75 Nm (wrist pitch). Three proprioceptive actuator types from Westwood Robotics were selected to meet these specifications, as shown in Table III.
@@ III-A3 Joint and Linkage
Since the elbow pitch actuator was relocated, a mechanism was required to transmit torque to the elbow joint. While timing belts offer simple power transmission, their low stiffness introduces compliance that limits torque control bandwidth [15]. Therefore, a dual 4-bar linkage mechanism in parallelogram configuration was implemented as shown in Fig. 2.
@@ III-A4 Tool Changing System
A manipulator with a general-purpose end-effector could be designed to handle several tasks, but would eventually limit the scope of the system. However, a manipulator that can utilize multiple specialized end-effectors is capable of being more effective in handling any number of tasks [16]. To allow YORI to better handle various tools for different tasks, YORI utilizes a pneumatic tool changer system. This allowed YORI to swap end-effectors between various cooking tools including a pan, a container, a knife, and a meat tenderizer. Each tool employs a custom tool plate, which the tool changer’s fingers can grip and lock onto using compressed air as shown in Fig. 3. A guiding cone feature allows the end-effector to correct any offset between the correct grasping position and the current position before the finger surface touches the tool plate. From our experiments, the tool changer could successfully guide itself to the locking position on the tool plate with up to 8 mm of translational offset from the center and with up to 10 degrees of angular offset.
@@ III-C Custom Tools and Appliances
Full integration and networking of cooking appliances was essential for system optimization. Rather than having the manipulator directly operate analog controls, which would introduce multiple failure points, an Internet of Things (IoT) approach was adopted [17]. Appliance selection was guided by four criteria: safety, robotic operability, cleanability, and multifunctionality. Based on common cooking operations, appliances were either modified from commercial units or custom-designed, as detailed in Table VI.
@@ III-C1 Rotating Mixer
The rotating mixer addresses mixing and tumbling operations that would otherwise limit multitasking capabilities if performed directly by the manipulator [18]. As shown in Fig. 5(c), the design features two rotation axes for varied movement patterns: fixed angles for stir-frying and swinging motions for mixing. The modular drum system enables quick tool changes and simplified cleaning.
@@ III-C2 Food Processor
The food processor employs a modular design addressing safety concerns associated with exposed blades [19]. As illustrated in Fig. 5(b), removable dicing chambers and blade assemblies facilitate cleaning and functional flexibility. During operation, pre-loaded ingredients in the dicing chamber are extruded through blades via a linear actuator-driven lid, with precise guidance from constraining rails. An electromagnet locks the chamber in place, while the manipulator positions a collection vessel below. This modularity supports various blade configurations for dicing, slicing, and pureeing operations.
@@ IV-E Task Scheduling
The system models cooking operations as a job-shop scheduling problem using Google OR-Tools’ CP-SAT solver [20]. Beyond traditional JSSP constraints, the formulation incorporates cooking-specific requirements.

### v4

# Yummy Operations Robot Initiative: An Autonomous Cooking System Utilizing a Modular Robotic Kitchen and Dual-Arm Proprioceptive Manipulator

@@ II-A Dual-Arm Manipulator
Cobots emerged in the late 1990s, enabling safe human-robot collaboration in tasks such as organizing, assembly and inspection. While lighter and more responsive than traditional industrial robots, cobots still suffer from relatively slow acceleration and high stiffness that limit performance in dynamic environments, and their weight remains inefficient for mobile platforms. These limitations have driven research toward lighter, backdrivable, and compliant manipulators with improved force control, exemplified by Berkeley’s Quasi-Direct Drive (QDD) BLUE [gealy2019quasi] and KoreaTech’s tendon-driven LIMS [kim2017anthropomorphic].
@@ II-B Proprioceptive Actuator
To enable backdrivability and compliance in manipulators, proprioceptive actuators, also known as QDD, offer an efficient solution. These actuators combine high-torque brushless DC motors with low reduction gearboxes (typically ¡10:1) to maintain torque transparency [Proprioceptive1]. Unlike traditional high-geared servo motors that suffer from high reflected inertia and poor compliance, proprioceptive actuators excel at backdrivability and impact mitigation [Proprioceptive2]. Their torque transparency enables direct torque estimation from motor current, eliminating the need for external force/torque sensors and simplifying both hardware and control systems.
@@ II-C Robotic Cell Layout Design
Robotic cell layout design extends traditional manufacturing layout optimization principles while addressing unique constraints of robotic systems, particularly in multi-robot environments and human-robot interaction scenarios [zhang2017challenges]. Recent advances in computational power have enabled sophisticated optimization techniques including Mixed Integer Linear Programming, Genetic Algorithms, and Convex Optimization to solve complex workspace design problems, with simulation-based optimization and digital twin technologies allowing virtual validation before physical implementation.
@@ II-D Manipulator Control
Accurate modeling is crucial for controlling manipulators with proprioceptive actuators, enabling precise torque calculation and force feedback for stable operation in dynamic environments. This accurate modeling enables two key control strategies: inverse dynamics control for trajectory tracking [lynch2017modern] and impedance control for safe interaction tasks [impedance]
@@ II-E Motion Planning
Minimizing jerk (time derivative of acceleration) is essential for smooth trajectory generation. Flash and Hogan showed human arm movements naturally minimize jerk using $L_{2}$ norm optimization [hogan1985], while Yazdani demonstrated that $L_{\infty}$ norm minimization yields biologically plausible ”bang-bang” control policies [cvxopt_min_jerk]. Although sampling-based methods with time parameterization are computationally efficient in confined environments, they often sacrifice smoothness and energy efficiency, leading to continued research in optimization-based techniques with jerk constraints [10610437]. Our work leverages YORI’s structured kitchen environment to eliminate dynamic obstacles, enabling full exploitation of minimum jerk optimization for smooth, energy-efficient dual-arm trajectories.
@@ II-F Task Scheduling
Effective task scheduling and real-time adaptability are crucial for robust automated cooking systems. The kitchen can be modeled as a finite state machine of heterogeneous robotic appliances, with orders represented as finite sequences containing operational parameters like cooking duration and seasoning amounts. These constraints, formulated using linear temporal logic (LTL) and computation tree logic (CTL), naturally align with job-shop scheduling problems (JSSPs). Prior work has applied dispatching rules [KitchenScheduling] and online MIP formulations [yi2022online] to cooking JSSPs. Building on these foundations, our work incorporates dynamic rescheduling logic and adaptive programming strategies to better handle real-world complexities including delays, failures, and continuous order arrivals.
@@ III Hardware Design
During the early stages of the project, we developed a prototype dual-arm manipulator using servo actuators to verify the feasibility of performing cooking tasks using robotic manipulators [YORI1]. However, due to the usage of rigid servo actuators and lack of force/torque sensors, the robot manipulator was unable to complete certain tasks where a precise torque estimation or rapid movement was required, such as cutting or tossing food. Furthermore, since all the components including actuators and structural parts are off-the-shelf products without any design optimizations, the payload at the end-effector was limited and the overall workspace was not ideal. With the numerous constructive experiences from the servo motor actuated manipulator prototype, we developed the next-generation proprioceptive dual-arm manipulator using high-performance quasi-direct drive actuators. These actuators allow for the manipulators compatibility with many cooking tasks, in particular those requiring impacts and precise force control, while being compliant to changes in the constructed environment.
@@ III-A Dual-Arm Manipulator
An initial servo-actuated prototype validated the joint configuration and link dimensions for cooking tasks [YORI1]. Building on this verified kinematic design, a next-generation dual-arm manipulator was developed using proprioceptive actuators, enabling impact-resilient manipulation with precise force control and environmental compliance, which are capabilities essential for dynamic cooking operations such as cutting and tossing.
@@ III-A1 Hardware Configuration
Link dimensions were determined based on human upper body proportions [human_measure] for compatibility with existing kitchen tools: upper arm (0.400 m), forearm (0.375 m), and clavicle (0.420 m) lengths allow both coordinated dual-arm tasks and independent operation. Carbon fiber tubes were selected for all arm links due to their optimal strength-to-weight ratio. The 1.3 m platform height ensures full workspace coverage. Table II summarizes the mechanical parameters.
@@ III-A2 Actuator Selection
Joint torque requirements were determined through static moment analysis for worst-case loading conditions (fully extended horizontal configuration) [YORI1]. Based on cooking task analysis, a 3 kg payload capacity was specified with a 1.5 safety factor, yielding maximum torque requirements of 37.93 Nm (shoulder pitch), 20.29 Nm (elbow pitch), and 3.75 Nm (wrist pitch). Three proprioceptive actuator types from Westwood Robotics were selected to meet these specifications, as shown in Table III.
@@ III-A3 Joint and Linkage
Since the elbow pitch actuator was relocated, a mechanism was required to transmit torque to the elbow joint. While timing belts offer simple power transmission, their low stiffness introduces compliance that limits torque control bandwidth [katz]. Therefore, a dual 4-bar linkage mechanism in parallelogram configuration was implemented as shown in Fig. 2.
@@ III-A4 Tool Changing System
A manipulator with a general-purpose end-effector could be designed to handle several tasks, but would eventually limit the scope of the system. However, a manipulator that can utilize multiple specialized end-effectors is capable of being more effective in handling any number of tasks [toolchanger]. To allow YORI to better handle various tools for different tasks, YORI utilizes a pneumatic tool changer system. This allowed YORI to swap end-effectors between various cooking tools including a pan, a container, a knife, and a meat tenderizer. Each tool employs a custom tool plate, which the tool changer’s fingers can grip and lock onto using compressed air as shown in Fig. 3. A guiding cone feature allows the end-effector to correct any offset between the correct grasping position and the current position before the finger surface touches the tool plate. From our experiments, the tool changer could successfully guide itself to the locking position on the tool plate with up to 8 mm of translational offset from the center and with up to 10 degrees of angular offset.
@@ III-C Custom Tools and Appliances
Full integration and networking of cooking appliances was essential for system optimization. Rather than having the manipulator directly operate analog controls, which would introduce multiple failure points, an Internet of Things (IoT) approach was adopted [IOT_Review]. Appliance selection was guided by four criteria: safety, robotic operability, cleanability, and multifunctionality. Based on common cooking operations, appliances were either modified from commercial units or custom-designed, as detailed in Table VI.
@@ III-C1 Rotating Mixer
The rotating mixer addresses mixing and tumbling operations that would otherwise limit multitasking capabilities if performed directly by the manipulator [Stir_Fry_arms]. As shown in Fig. 5(c), the design features two rotation axes for varied movement patterns: fixed angles for stir-frying and swinging motions for mixing. The modular drum system enables quick tool changes and simplified cleaning.
@@ III-C2 Food Processor
The food processor employs a modular design addressing safety concerns associated with exposed blades [Robot_Cutting]. As illustrated in Fig. 5(b), removable dicing chambers and blade assemblies facilitate cleaning and functional flexibility. During operation, pre-loaded ingredients in the dicing chamber are extruded through blades via a linear actuator-driven lid, with precise guidance from constraining rails. An electromagnet locks the chamber in place, while the manipulator positions a collection vessel below. This modularity supports various blade configurations for dicing, slicing, and pureeing operations.
@@ IV-E Task Scheduling
The system models cooking operations as a job-shop scheduling problem using Google OR-Tools’ CP-SAT solver [cpsatlp]. Beyond traditional JSSP constraints, the formulation incorporates cooking-specific requirements.
@@ Acknowledgments
Donghun Noh Robotics and Mechanisms Laboratory, the Department of Mechanical and Aerospace Engineering, University of California, Los Angeles, CA 90095, USA. Email: donghun.noh@ucla.edu
Hyunwoo Nam Robotics and Mechanisms Laboratory, the Department of Mechanical and Aerospace Engineering, University of California, Los Angeles, CA 90095, USA. Email: harrynam@ucla.edu
Kyle Gillespie Robotics and Mechanisms Laboratory, the Department of Mechanical and Aerospace Engineering, University of California, Los Angeles, CA 90095, USA. Email: ktgillespie@ucla.edu
Yeting Liu Robotics and Mechanisms Laboratory, the Department of Mechanical and Aerospace Engineering, University of California, Los Angeles, CA 90095, USA. Email: liu1995@ucla.edu
Dennis Hong Robotics and Mechanisms Laboratory, the Department of Mechanical and Aerospace Engineering, University of California, Los Angeles, CA 90095, USA. Email: dennishong@ucla.edu